### v2

# Performance of a coarsely pixelated LAPPD photosensor for the SoLID light gas Cherenkov detector

Abstract
The SoLID spectrometer light gas Cherenkov counter requires photosensors that operate in a high luminosity and high background environment. In its reference design, a configuration of 9 or 16 tiled multi-anode photomultipliers (MaPMTs) covers the sensing area. To evaluate the use of a pixelated INCOM Large Area Picosecond Photodetector (LAPPD) as an alternative to cover the same area, we tested the performance of this photosensor in realistic SoLID running conditions in Hall C at the Thomas Jefferson National Accelerator Facility (TJNAF, a.k.a Jefferson Lab or JLab). The results of this test confirmed that the coarse-pixelated (2.5x2.5 cm pixel size) LAPPD is capable of handling the total projected signal and background rates of the three pillar experiments of SoLID. The photosensor detected Cherenkov signals with the capability of separating single-electron events from pair production events while rejecting background. Although the design was not aimed at Ring Imaging Cherenkovs, Cherenkov disk images were captured in two gas radiators. Through a direct comparison with a GEANT4 simulation, we confirmed the experimental performance of the LAPPD.
1 Introduction
The Solenoidal Large Intensity Device (SoLID) [1, 2] is a new spectrometer designed to exploit the full potential of the 12 GeV energy upgrade of the Continuous Electron Beam Accelerator Facility (CEBAF) at Jefferson Lab (JLab). The SoLID spectrometer will be operated at the full CEBAF luminosity of 10 cms [2] in the Parity Violation Deep Inelastic Scattering (PVDIS) experiment and at a reduced luminosity of 10 cms in the Semi-Inclusive Deep Inelastic Scattering (SIDIS) or Exclusive production near threshold, known as the SIDIS-J/psi experiments. In the SIDIS- SoLID spectrometer configuration, Cherenkov counters were designed for a trigger system and particle identification: a light gas Cherenkov (LGC) for e/ separation and a heavy gas Cherenkov (HGC) for /K and /p separation. Hamamatsu flat panel multi-anode photomultipliers (MaPMTs) [3] were selected as the baseline photosensors. The MaPMTs have 64 pixels with sensitivity down to single photon detection. Signals from pixels can be either read out individually (64 pixels) or as 16 quad-sum channels. However, a major drawback of this sensor is the loss of efficiency in a magnetic field [4], and thus magnetic shielding is unavoidable for it to be used in a magnetic field. An alternative sensor is the Microchannel plate photomultiplier (MCP-PMT), a photosensor with a proven strong magnetic field resiliency, and high granularity. Its design consists of a stacked pair of MCPs in a chevron configuration for secondary electron amplification. Due to the significantly reduced electron amplification path, the MCP-PMT shows an exceptional magnetic field tolerance [5], a great advantage to avoid heavy and cumbersome mu-metal magnetic shielding in large array configurations of MAPMTs in each counter sector of SoLID. Unfortunately, the high cost of these commercial MCP-PMTs prevents us from readily tiling them for large area coverage Cherenkov counters in the presence of strong magnetic fields.
Over the past decade a new type of MCP-PMT, the large area picosecond photodetector (LAPPD) was developed by the LAPPD collaboration [5] and more recently commercialized by Incom, Inc [7, 8, 9]. The LAPPDs apply resistive and secondary emissive layers on a large area of glass capillary through atomic layer deposition (ALD), achieving low-cost functional MCPs, thus the final cost of the manufactured device is expected to be much lower per active area than that of traditional MCP-PMTs. Meanwhile, the performance of the LAPPD is comparable to that of the standard commercial MCP-PMTs, including timing resolution, position resolution, and dark count. The expected low cost and high performance make the LAPPD an ideal photosensor candidate for Cherenkov counters.
To explore the potential of using a LAPPD as the photosensor for the SoLID Cherenkov counter and evaluate its performance in a high-rate background environment, we performed an earlier opportunistic but limited test [10] of a strip-line read-out Gen-I LAPPD at Jefferson Lab. Although the event separation was not very clean due to a low quantum efficiency and non-pixelated readout the results showed a promising performance of the GEN-I LAPPD in a harsh running environment of Hall C at Jefferson Lab. With the availability of Gen-II (capacitive coupling) pixelated LAPPD, a more comprehensive test was designed and performed to validate it as a viable photosensor for the SoLID gas Cherenkov counters as part of a pre-R&D. In this paper, we describe the details of this comprehensive test, with the experimental setup, the Gen-II LAPPD geometry and data taking procedure. Results will be presented and discussed as we evaluate the potential of a coarse-pixelated LAPPD as a photosensor candidate for the SoLID gas Cherenkov counters.
2 Telescopic Cherenkov device design
A Telescopic Cherenkov Device (TCD) LAPPD was designed and constructed, as shown in Figure 1. Compared to Ref. [10] the present design to allow the use of an array of 4x4 MaPMTs or a 20x20 cm LAPPD. It consists of the primary Cherenkov tank, the LAPPD electronics housing box, the aluminum-coated reflective mirror, aluminum foil end caps, the gas supply system, the power supply system, and readout electronics.
The primary Cherenkov tank was assembled from 14" diameter PVC pipes, and machine cut to the required dimensions. Inside the tank, felt flocking was attached to minimize secondary reflections. Additional apertures were added to allow valves and gas flow, sensors, and an LED test light source. A special housing box was designed to accommodate the coarse-pixelated LAPPD and its electronic cables. The housing box was designed to maintain structural and hermetic integrity while allowing detachable access plates in order to reach the enclosed electronics. To secure the removable LAPPD electronics housing box, large circular rubber gaskets were applied. Aluminum windows and PVC window frames were pressure tested to ensure the system was airtight. Flat elliptical mirrors were constructed from carbon-fiber bases and Lexan reflective film. The primary mirror was mounted inside the tank and aligned to reflect light incident down the central ray of the tank and reflected by the mirror onto the center of the photon detector array. The LAPPD housing box was surrounded with lead bricks during beamtime to protect the photosensor from radiation damage. A 370 nm light-emitting device (LED) was installed near the entrance window for calibration and checking purposes. A pressure regulation system was assembled, it included a gauge pressure manometer, two solenoid valves, a desktop PID controller with a laptop interface, and the required cabling. Operating at atmospheric pressure helped simplify the mechanical design and minimized the window thicknesses. Carbon dioxide (CO) and octafluorocyclobutane (CF) were used as the radiator medium in this system to evaluate the LAPPD performance for both light gas Cherenkov and heavy gas Cherenkov (see next section). In order to ensure high gas purity and before stabilizing it at atmospheric pressure, the telescopic Cherenkov device was flushed in several cycles each time a new gas was used.
3 Large area picosecond photodetector (LAPPD)
The LAPPD photodetector is a novel microchannel plate photomultiplier based on the use of a newly developed microchannel plate, functionalized through atomic layer deposition on a glass capillary array structure (GCA-ALD-MCP) [7]. Figure 2(a) shows an exploded-view drawing of the LAPPD, consisting of a top window with an alkali photocathode on the inner surface to generate photoelectrons once incident photons enter the window, two microchannel plates with Chevron configuration to amplify the number of secondary electrons and a bottom plate to collect the resulting charge cluster for readout.
The tested Gen-II LAPPD features a 64-pixelated readout with a pixel size of 25 mm x 25 mm through capacitive coupling. The entrance window of this LAPPD is regular B33 glass which absorbs photons below 270 nm. To enhance the Cherenkov photons detection efficiency a lower wavelengths, a thin layer of p-terphenyl wavelength shifter coating was applied [11] to convert ultraviolet photons into visible photons for detection. Fig.1(a) shows the exploded-view drawing of LAPPD and an image of the tested Gen-II coarse-pixelated LAPPD installed in the telescopic Cherenkov housing box. The detailed parameters and performance on the bench of this particular LAPPD are listed in Table 1.
| Incom Gen-II LAPPD | |
| PMT type | MCP-PMT |
| Device size | 20 20 cm |
| Active area | 20 20 cm |
| Readout type | Pixel (2.5cm 2.5cm) |
| Entrance window material | B33 glass |
| Wavelength shifter coating | Yes |
| Response wavelength range | 185 – 650 nm |
| Quantum efficiency (QE) | 15% (17% maximum) |
| Operation high voltage | 2350 V (MCP = 875 V) |
| Gain | 9.5 × 10 |
4 Experimental setup
The telescopic Cherenkov device was installed in hall C at Jefferson Lab, along with the Super High Momentum Spectrometer (SHMS). The experiment was performed parasitically with the ongoing Hall C fixed-target experiments, the setup location was chosen for minimal interference with the running experiment as well as to have a moderate rate background environment. The Cherenkov was positioned to face the entrance window of the liquid hydrogen target to allow the scattered electrons from the target to freely pass through the entrance window after crossing the target and the target scattering chamber aluminum walls.
The experiment used an 8.1 GeV electron beam, impinging a liquid H target. While the scattered electrons or produced particles pass through the tank, Cherenkov light is produced along the particle path if its velocity fulfills the Cherenkov light emission condition. The Cherenkov light is then reflected at 90 by the mirror at the section, and collected by the LAPPD. The coarse-pixelated LAPPD was accommodated in the housing box, and surrounded by lead blocks to prevent radiation damage. Scintillator paddles and Shashlik electromagnetic calorimeter blocks [12] were placed in front of the entrance window and behind the exit window respectively to generate coincident external triggers. All triggers and signals from the scintillator paddles, the calorimeter, and LAPPD were fanned out and sent to the F250 flash Analog-to-Digital Converters (FADC250) [13] for data acquisition.
The current SoLID detector design includes two Cherenkov counter sub-systems: the light gas Cherenkov (LGC) and the heavy gas Cherenkov (HGC). The light gas Cherenkov applies CO at 1 atm (index of refraction n = 1.00045) as a gas radiator, it will be mainly used to separate electrons from pions giving a very low momentum threshold for electrons (0.017 GeV) and a medium momentum threshold for pions (4.75 GeV). The heavy gas Cherenkov applies CF at 1.7 atm (index of refraction n = xxx) as gas radiator, it will provide a clear distinction between charged pion and kaon Cherenkov signals from a momentum range of 2.5 GeV/c to 7.5 GeV/c for the SoLID SIDIS program at forward angles from 8 to 15. To evaluate the performance of pixelated LAPPD in both the light gas and heavy gas Cherenkov sub-systems, we experimented with CO and CF as the gas radiators, respectively.
The Cherenkov counter was installed in Hall C at Jefferson Lab, as shown in Figure 3. It was mounted at beamline level pointing towards the target cell and positioned 4.8 m away from the target scattering chamber. Its entrance window was facing the target at an angle of 68 with respect to the beam direction. The distance and angle were constrained by space and safety considerations due to the ongoing experiment in the hall, resulting in a moderate background-rate environment. Lead blocks shielded the photosensors from direct radiation damage.
5 Simulation
In order to understand the experimental test results and later optimize the SoLID LGC design, we conducted a thorough simulation of the entire Cherenkov process. A Geant4 simulation [14] of the TCD, including the Cherenkov counter, the EM calorimeter (EC), and scintillator (SC) detectors, was developed similar to the SoLID simulation. The setup is shown in Figure 4. The measured LAPPD quantum efficiency was used and the Cherenkov optical physics cut-off was chosen to be 200 nm. We studied the resulting number of photo-electrons in the Cherenkov detector from two main sources, namely high energy scattered electrons and /0 production with subsequent decay photons conversion into electron-positron pairs from the polarized He gas target. Simulation results are presented together with the experimental test results for comparison in the following sections.
The TCD setup in the Geant4 simulation.
6 Results and Discussion
The LAPPD has a signal readout scheme of 8 × 8 array of 25 mm × 25 mm pixels. Signals from all 64 channels were recorded during the experiment, and the data was analyzed offline. In a Cherenkov signal event, the generated photons will be distributed as a light disk at the LAPPD photosensor. Multiple pixels should collect Cherenkov photons and fire simultaneously. Figure 5 shows the signal sum for different numbers of pixels (N) subgroups. The signals can be categorized into three groups: i) the random coincidence of single photo-electron signals with N = 1, 2, 3, demonstrating a significant suppression when requiring multiple pixels (channels) to fire within the same time window. ii) Cherenkov signals with N = 9, 10, 11. The signals follow a Gaussian distribution centered around the number of photoelectrons (NPE) of 17, corresponding to central hits. iii) Gaussian-like signals with N = 19, 20, 21, approximately double the NPE of the typical Cherenkov events. These high-NPE signals are possibly the coincidence sums of simultaneous detection of an electron and positron from a pair production process.
The LAPPD test includes different types of radiator gas: CO, a mixture of 80% CF and 20% CO, and CF. The normalized distributions of the signal sum from these types of gases are shown in Figure 6. All three distributions show an identical peak below 100 ADC channel, indicating a stable random coincidence background from the high rate environment during the beam test. The heavy gas (CF) shows a more prominent bump from 250 to 1000 ADC channels, indicating more photons detected from typical Cherenkov events.
An important feature of the Cherenkov counter is the detection of the photon disk with photosensors. The disk size is related to the particle’s momentum, the refraction index of the radiator, and the photon’s path length to the photosensors. With the light gas radiator (CO), a typical Cherenkov photon ring would have 9 to 10 fired pixels. In heavy gas (CF), the Cherenkov light cone contains more photons and has a larger opening angle, resulting in a greater disk radius on the detection plane with more fired pixels. According to their refractive index (n = 1.0005 and n = 1.0014), the Cherenkov disk from CF is about 70% larger in radius than the disk from CO. Thus, the corresponding coverage of a typical Cherenkov signal is 15 to 16 pixels in the heavy gas. The top panel of Figure 7 shows the Cherenkov rings we observed with the LAPPD and different gas radiators from the large scattering angle beam test. The event samples show a small shift of the signal center to the right of the geometrical center and a slightly wider distribution, compared to the simulation shown in the bottom panel of Figure 7. Without a tracking detector, we cannot precisely select the events that enter the telescope Cherenkov detector right on its geometrical center, so shifts from event to event are expected. Most importantly, the size and shape of the Chereknov rings observed from the data are consistent with those in the simulation.
Given the expected light disk sizes for typical Cherenkov events in different gas radiators, we sum all the signals with the requirement of 9 to 10 fired pixels in CO and 15 to 16 in CF. The signal sums with the LAPPD sensor show an NPE = 12 in CO and NPE = 21 in CF. Our simulation can also well describe the difference 20 between the two gas radiators. The simulation yields NPE = 9 in CO and NPE = 23 in C4F8, showing a good agreement with the data. The results are presented in Figure 8.
7 Summary
We reported on the performance study of a coarse-pixelated LAPPD using a telescopic Cherenkov device in an open high-rate environment at Jefferson Lab. A pixelated Gen-II LAPPD was tested with two different radiators, a light gas CO and a heavy gas CF that will be used in the SoLID experiment. The coarse-pixelated Gen-II LAPPD device was able to operate in such an environment for both gas radiators, detect the Cherenkov signals, and separate single-electron events and pair production events. With this LAPPD a typical Cherenkov disk was observed in our TCD, with a smaller diameter for CO radiator and a larger diameter for CF radiator. A larger number of photoelectrons was generated from the signal sum in CF compared to CO case as expected. GEANT 4 simulations were performed to confirm the results, showing a good agreement with the data. Our experiment confirms that for SoLID the pixelated Gen-II LAPPD could handle the high background rate environment for the more stringent requirement of the SoLID PVDIS experiment.
Acknowledgement
Argonne National Laboratory’s work was supported by the US Department of Energy, Office of Science, Office of Nuclear Physics, under contract DE-AC02-06CH11357. Part of the work was supported by Laboratory Directed Research and Development (LDRD) funding from Argonne National Laboratory, provided by the Director, Office of Science, of the US Department of Energy under contract DE-AC02-06CH11357. The work at Temple University is supported by the US Department of Energy award DE-SC0016577. Incom’s work is supported by the U. S. Department of Energy, Office of Science, Office of Basic Energy Sciences, Offices of High Energy Physics and Nuclear Physics under DOE contracts: DE-SC0015267, DE-SC0017929, DE-SC0018778, and DE-SC0019821. This material is based upon work supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under contract DE-AC05-06OR23177.